# Debiasing Welch’s Method for Spectral Density Estimation

### 21 Dec 2023 v1

@@ Abstract
Welch’s method provides an estimator of the power spectral density that is statistically consistent. This is achieved by averaging over periodograms calculated from overlapping segments of a time series. For a finite length time series, while the variance of the estimator decreases as the number of segments increase, the magnitude of the estimator’s bias increases: a bias-variance trade-off ensues when setting the segment number. We address this issue by providing a a novel method for debiasing Welch’s method which maintains the computational complexity and asymptotic consistency, and leads to improved finite-sample performance. Theoretical results are given for fourth-order stationary processes with finite fourth-order moments and absolutely continuous fourth-order cumulant spectrum. The significant bias reduction is demonstrated with numerical simulation and an application to real-world data, where several empirical metrics indicate our debiased estimator compares favourably to Welch’s. Our estimator also permits irregular spacing over frequency and we demonstrate how this may be employed for signal compression and further variance reduction. Code accompanying this work is available in the R and python languages.
Keywords: Welch’s method; bias reduction; tapering; non-parametric estimation; periodogram. ${}^{1}$ School of Physics, Mathematics & Computing, The University of Western Australia, 35 Stirling Hwy, Crawley, Australia ${}^{2}$ Oceans’ Graduate School, The University of Western Australia, 35 Stirling Hwy, Crawley, Australia ${}^{3}$ Department of Mathematics, Imperial College London, South Kensington, London SW7 2BX, U.K.
@@ 1 Introduction
The periodogram is the fundamental non-parametric estimator of the power spectral density; however, it suffers from two deficiencies: the periodogram is statistically inconsistent and is biased with finite samples. By averaging multiple periodograms calculated from partitioned or overlapping segments of a time series, Welch (1967) overcomes the issue of inconsistency to provide a statistically consistent estimator of the power spectral density, eponymously known as Welch’s estimator. Welch’s estimator is simple to implement and easy to understand: all reasons that has led to its ubiquity in the applied sciences. However, for finite length time series, the issue of bias remains. Each segment’s periodogram is biased and Welch’s estimator inherits this bias. Further, as the magnitude of the bias is inversely related to the length of the segments, there is the classical inverse relationship between bias and variance of the estimator; this leads to the undesirable property that we become increasingly more certain in an estimator that is increasingly more biased. In this paper, we introduce a debiased analogue to Welch’s estimator that we entitle the debiased Welch estimator.
Historically, a number of methodologies have been developed to mitigate the effects of bias. One such methodology is tapering (Thomson, 1982), where a taper is applied to the data; although, typically this reduces correlation between distant frequencies at the cost of increasing correlation between nearby frequencies, and so bias remains. More recently, Sykulski et al. (2019) provided methodology to debias spectral density estimates, but this methodology relies on the availability of the parametric form of the power spectral density. In this paper, we extend the results in Sykulski et al. (2019) to non-parameteric estimation by constructing a natural basis representation to the power spectral density. From this basis representation we intentionally bias the bases, fit the biased bases to a biased estimate of the power spectral density, and recover debiased coefficients that are used to form our debiased non-parametric estimator. Tapering is ubiquitous in the applied sciences and augments both parametric and non-parametric estimation of the power spectral density, hence our method is developed to be applicable with any choice of taper.
@@
Novel insight is provided across two theorems. The first presents theoretical results that make explicit how the properties of Welch’s estimator are a function of the number of observations, number of segments, and the segment lengths. This theorem makes the finite sample bias explicit and motivates the construction of a new debiased alternative which we propose via a basis representation. To obtain a good representation of the power spectral density, however, we must specify a large number of basis functions and so estimation of the coefficients via direct optimisation of the Whittle likelihood, as in Sykulski et al. (2019), becomes computationally infeasible. The major contribution of this work is therefore housed in the second theorem, where we introduce methodology that leverages results established in the first theorem, to provide asymptotic results that enable fast and direct computation of the debiased Welch estimator. We build theory for time series requiring fourth-order stationarity with finite fourth-order moments and absolutely continuous fourth-order cumulant spectrum. The theory is established for all tapers, subject to the standard conditions of squared summability of the tapering sequence, thereby extending many classical results that use the raw periodogram, for example, those in Anderson (1971), Brillinger (1975) and Katznelson (1968).
@@
The theoretical developments are also demonstrated by simulation. First, we use a canonical auto-regressive model of order four studied throughout the spectral estimation literature, see for example Percival et al. (1993). We use this model to contrast the performance of the standard and debiased Welch estimators as a function of the number of segments and the segment length. Next, we demonstrate the debiased Welch estimator in a real-world case study of coastal wave measurements. Here, we provide an example of where parametric methods struggle to capture the complexity of the process, thus necessitating the use of non-parametric methods. Finally, unlike Welch’s estimator, the debiased Welch estimator does not require even spacing over frequency; we use this to demonstrate how we may simultaneously debias and compress the signal. We show this for a simulation of a stochastic process with log-linear decay in the power spectral density over frequency, a pertinent case-study for many applications. Code accompanies our results on multiple software platforms; both the R and python implementations are available at github.com/TIDE-ITRH.
@@ 2.1 Fundamentals and definitions
Define by $\{X_{t}\}$ a discrete real-valued stochastic process, sampled at interval $\Delta$ and indexed by $t\in\mathcal{Z}$, where $\mathcal{Z}$ is the set of integers. The process $\{X_{t}\}$ may be either thought of as a true discrete-time stochastic process, such as an auto-regressive model, or as uniformly sampled observations of a continuous-time stochastic process $\{X(t)\}$ in which case $X_{t}=X(\Delta t)$. We assume fourth-order stationarity, with finite fourth-order moments and absolutely continuous fourth-order cumulant spectrum, $\kappa(\alpha,\beta,\zeta)$. Define the auto-covariance sequence of $\{X_{t}\}$ as $\gamma(\tau)=\mathrm{E}[X_{t}X_{t-\tau}]$ for $\tau\in\mathcal{Z}$. Following Wiener–Khinchin’s theorem, a necessary and sufficient condition for $\gamma(\tau)$ to be the auto-covariance function of $\{X_{t}\}$ is that there exists a power spectral distribution function $F(\omega)$ such that
@@
for $\omega\in[-\pi/\Delta,\pi/\Delta]$, where (1) is a Riemann-Stieltjes integral and $\omega$ is defined in radians. Absolute continuity of $\kappa(\alpha,\beta,\zeta)$ implies absolute summability in $\gamma(\tau)$. Consequently, there exists a uniformly continuous power spectral density (see, for instance, Section 4.2 Shumway and Stoffer, 2000), $f(\omega)$, such that $f(\omega)\mathrm{d}\omega=\mathrm{d}F(\omega)$, and $f(\omega)$ forms a Fourier pair with $\gamma(\tau)$ so that
@@ 2.3 Welch’s method for spectral density estimation
where $S$ is an integer-valued shift factor satisfying $0<S\leq L$, so that $p=(L-S)/L$ is the percentage overlap of the blocks, and the relationship between $M$, $L$ and $p$ is
$ML(1-p)+Lp=n,$ (8)
@@
where $\rho(\omega;|m-m^{\prime}|,h)=\mathrm{cor}\left[I_{L}^{m}(\omega;h),I_{L}^{m^{\prime}}(\omega;h)\right]$ and $\mathrm{var}[I_{L}(\omega;h)]=\mathrm{var}[I_{L}^{m}(\omega;h)]$ for all $m$. We establish some important properties of Welch’s estimator in Theorem 1.
@@
Assume a fourth-order stationary stochastic process $\{X_{t}\}$ with finite fourth-order moments and absolutely continuous fourth-order cumulant spectrum. Welch’s estimator, defined in (9), of the observed process $\mathrm{X}_{n}=(x_{0},\dots,x_{n-1})$, with data taper $\{h_{t}\}$ that satisfies $\sum h_{t}^{2}=1$, has the following properties.
@@
Proof. Proofs are provided in the Supplementary Material.
@@
As $n$ increases, then it is not possible to increase both $L$, the block length, and $M$, the number of blocks, linearly with $n$ for a fixed overlap $p$, see (8); a trade-off is required. For fixed $L$ and linearly increasing $M$, Theorem 1a establishes that Welch’s estimate has the optimal convergence rate of $\mathcal{O}_{p}(M^{-1/2})=\mathcal{O}_{p}(n^{-1/2})$, but the estimator is inconsistent for all processes except white noise due to the bias inherent in $f(\omega)*\mathcal{H}_{L}(\omega)$ for fixed $L$ as discussed. On the other extreme, for fixed $M$ and linearly increasing $L$, the bias asymptotically vanishes but the variance is $\mathcal{O}_{p}(1)$ (Theorem 1a) and therefore the Welch estimate suffers the same inconsistency as the periodogram (Theorem 1b). Consistency is only achieved by increasing both $L$ and $M$ with $n$ at reduced rates.
@@ 3.3 Model fitting
The motivation for (16) is as follows. From Theorem 1b, as $\lim_{L\rightarrow\infty}\mathrm{var}\left[\bar{I}_{L}(\omega;h)\right]=c\;\mathrm{var}\left[I_{L}(\omega;h)\right]$, over all $\omega$, we substitute $\mathrm{var}\left[I_{L}(\omega;h)\right]$ for $\mathrm{var}\left[\bar{I}_{L}(\omega;h)\right]$ in (15). Further, as we establish in Theorem 2 below, $\bar{I}_{L}(\omega;h)^{2}$ converges in $M$ and $L$ to $\mathrm{var}[I_{L}(\omega;h)]$, and so we approximate $\mathrm{var}[I_{L}(\omega;h)]$ by $\bar{I}_{L}(\omega;h)^{2}$, resulting in (16). From $\hat{\vartheta}=(\hat{a}_{1},\dots,\hat{a}_{K})$, we define a debiased functional estimate of $f(\omega)$ as $f(\omega;\hat{\vartheta})=\sum_{k=1}^{K}\hat{a}_{k}B_{k}(\omega)$. For the symmetric rectangular basis function of Section 3.2, we define the debiased Welch estimator at a discrete set of frequencies as $\hat{I}(\omega_{k})=\hat{a}_{k}B_{k}(\omega_{k})=\hat{a}_{k}/\delta_{k}$ for $\omega_{k}\in\{\omega_{1}^{\mathrm{c}},\dots,\omega_{K}^{\mathrm{c}}\}$; this is analogous to the frequencies for which Welch’s estimator is generally defined.
@@
Assume a fourth-order stationary process, $\{X_{t}\}$, with finite fourth-order moments and absolutely continuous fourth-order cumulant spectrum, $\kappa(\alpha,\beta,\zeta)$. The variance of the periodogram of any observed length-$L$ segment of $\mathrm{X}_{n}=(x_{0},\dots,x_{n-1})$, for all $\omega\neq 0,\omega_{\mathcal{N}}$, and with data taper $\{h_{t}\}$ that satisfies $\sum h_{t}^{2}=1$, is
@@
    $\mathrm{var}[I_{L}(\omega;h)]=\bar{I}_{L}(\omega;h)^{2}+\mathcal{O}_{p}\left(\frac{1}{\sqrt{M}}+\frac{\log^{2}L}{L^{2}}\right)$, when $\kappa(\alpha,\beta,\zeta)=0$ for all $\alpha,\beta,\zeta$; and
@@
Proof. The proof is provided in the Supplementary Material.
@@
From Theorem 1 of Sykulski et al. (2019), we find that the debiased Welch estimator, $\hat{I}(\omega_{k})$, when computed through (6), is unbiased so that $\mathrm{E}[\hat{I}(\omega_{k})]=f(\omega_{k};\vartheta)$. The weighted least squares routine in (16) follows asymptotic arguments in Theorems 1 and 2 that result in an estimator that is asymptotically unbiased and $\mathcal{O}_{p}(M^{-1/2})$ as long as $L$ and $M$ both increase with $n$, and significantly debiased for finite samples compared to the standard Welch estimator, as we will demonstrate in the following sections through simulation and applications.

### 11 Apr 2024 v2

@@ Abstract
Welch’s method provides an estimator of the power spectral density that is statistically consistent. This is achieved by averaging over periodograms calculated from overlapping segments of a time series. For a finite length time series, while the variance of the estimator decreases as the number of segments increase, the magnitude of the estimator’s bias increases: a bias-variance trade-off ensues when setting the segment number. We address this issue by providing a novel method for debiasing Welch’s method which maintains the computational complexity and asymptotic consistency, and leads to improved finite-sample performance. Theoretical results are given for fourth-order stationary processes with finite fourth-order moments and absolutely convergent fourth-order cumulant function. The significant bias reduction is demonstrated with numerical simulation and an application to real-world data. Our estimator also permits irregular spacing over frequency and we demonstrate how this may be employed for signal compression and further variance reduction. Code accompanying this work is available in R and python.
Keywords: Welch’s method; bias reduction; tapering; non-parametric estimation; periodogram. ${}^{1}$ School of Physics, Mathematics & Computing, The University of Western Australia, 35 Stirling Hwy, Crawley, Australia ${}^{2}$ Oceans’ Graduate School, The University of Western Australia, 35 Stirling Hwy, Crawley, Australia ${}^{3}$ Department of Mathematics, Imperial College London, South Kensington, London SW7 2AX, U.K.
@@ 1 Introduction
The periodogram is the fundamental non-parametric estimator of the power spectral density; however, it suffers from two deficiencies: the periodogram is statistically inconsistent and is biased with finite samples. By averaging multiple periodograms calculated from partitioned or overlapping segments of a time series, Welch (1967) overcomes the issue of inconsistency to provide a statistically consistent estimator of the power spectral density, eponymously known as Welch’s estimator. Welch’s estimator is simple to implement and easy to understand: all reasons that has led to its ubiquity in the applied sciences. However, for finite length time series, the issue of bias remains. Each segment’s periodogram is biased and Welch’s estimator inherits this bias. Further, as the magnitude of the bias is inversely related to the length of the segments, there is the classical inverse relationship between bias and variance of the estimator; this leads to the undesirable property that we become increasingly more certain in an estimator that is increasingly more biased. The bias is due to a phenomenon called blurring, which is observed most prominently in low power regions, and so it is often ignored as an unimportant aspect of the spectrum; although, such bias may severely impact inference and scientific understanding (e.g. Thomson et al., 1995). In this paper, we introduce a debiased analogue to Welch’s estimator that we entitle the debiased Welch estimator.
Welch (1967) popularised estimating a spectral density by averaging tapered periodograms calculated from shifted overlapping windows of a time series, thus generalising Bartlett’s estimator (Bartlett, 1950). Statistical properties of Welch’s estimator, and its behaviour under varying tapers, parametrisations, and smoothness assumptions of the underlying density, have been established; see, for instance, Nuttall (1971), Zhurbenko (1980, 1986), Dahlhaus (1985), and (Politis, 2005). Welch’s estimator may further been seen as a special case of modern methodologies such as scalable subbaging (Politis, 2024). In many cases, these developments improve upon the standard statistical properties of Welch’s estimator; however, none consider the explicit debiasing under general conditions. Alternative non-parametric methodologies have been developed to mitigate the effects of bias, most prominently lag-windowing (Blackman and Tukey, 1958) and multi-tapering (Thomson, 1982); although, typically these reduce correlation between distant frequencies at the cost of increasing correlation between nearby frequencies, and so bias remains. In this paper, we extend the results in Sykulski et al. (2019), who provided methodology to debias parametric spectral density estimates, to non-parametric estimation by constructing a natural basis representation to the power spectral density. From this basis representation we intentionally bias the bases, fit the biased bases to a biased estimate of the power spectral density, and recover debiased coefficients that are used to form our debiased non-parametric estimator.
@@
Novel insight is provided across two theorems. The first presents theoretical asymptotic properties of Welch’s estimator, making the finite sample bias explicit and motivating the construction of a new debiased alternative which we propose via a basis representation. To obtain a good representation of the power spectral density, we must specify a large number of basis functions and so estimation of the coefficients via direct optimisation of the Whittle likelihood, as in Sykulski et al. (2019), becomes computationally infeasible. The major contribution of this work is therefore housed in the second theorem, where we introduce methodology that leverages results established in the first theorem, to provide asymptotic results that enable fast and direct computation of the debiased Welch estimator. We build theory for time series requiring fourth-order stationarity with finite fourth-order moments and absolutely convergent fourth-order cumulant function. The theory is established for all tapers, subject to the standard conditions of squared summability of the tapering sequence, thereby extending many classical results that use the raw periodogram, for example, those in Anderson (1971), Brillinger (1975) and Katznelson (1968).
@@
The theoretical developments are also demonstrated by simulation. To evidence the asymptotics, we use a canonical auto-regressive model of order four studied throughout the spectral estimation literature (see for example Percival et al., 1993); then, we present results for shorter-length time series generated from a Matérn covariance function. Next, we demonstrate the debiased Welch estimator in a real-world case study of coastal wave measurements. This provides an example of where parametric methods struggle to capture the complexity of the process, thus necessitating the use of non-parametric methods. Finally, unlike Welch’s estimator, the debiased Welch estimator does not require even spacing over frequency; we use this to demonstrate how we may simultaneously debias and compress the signal. Code accompanies our results; R and python implementations are available at github.com/TIDE-ITRH.
@@ 2.1 Fundamentals and definitions
Define by $\{X_{t}\}$ a discrete real-valued stochastic process, sampled at interval $\Delta$ and indexed by $t\in\mathcal{Z}$, where $\mathcal{Z}$ is the set of integers. The process $\{X_{t}\}$ may be either thought of as a true discrete-time stochastic process, such as an auto-regressive model, or as uniformly sampled observations of a continuous-time stochastic process $\{X(t)\}$ in which case $X_{t}=X(\Delta t)$. Define the auto-covariance sequence of $\{X_{t}\}$ as $\gamma(\tau)=\mathrm{E}[X_{t}X_{t-\tau}]$ for $\tau\in\mathcal{Z}$. Following Wiener–Khinchin’s theorem, a necessary and sufficient condition for $\gamma(\tau)$ to be the auto-covariance function of $\{X_{t}\}$ is that there exists a power spectral distribution function $F(\omega)$ such that
@@
for $\omega\in[-\pi/\Delta,\pi/\Delta]$ defined in radians, where (1) is a Riemann-Stieltjes integral. We assume fourth-order stationarity, with finite fourth-order moments and fourth-order cumulant function
$\mathcal{K}(\tau_{i},\tau_{j},\tau_{k})=\mathrm{E}[X_{0}X_{\tau_{i}}X_{\tau_{j}}X_{\tau_{k}}]-\gamma(\tau_{i})\gamma(\tau_{j}-\tau_{k})-\gamma(\tau_{j})\gamma(\tau_{i}-\tau_{k})-\gamma(\tau_{k})\gamma(\tau_{i}-\tau_{j})$
that is absolutely convergent. As $\sum_{\tau_{i},\tau_{j}=-\infty}^{\infty}|\gamma(\tau_{i})\gamma(\tau_{j})|$ is convergent only if $\sum_{\tau=\infty}^{\infty}|\gamma(\tau)|<\infty$, then $\gamma(\tau)$ too is absolutely convergent. Consequently, the power spectral distribution is absolutely continuous (see Section 4.2 Shumway and Stoffer, 2000), and so there exists a power spectral density $f(\omega)$ that satisfies $f(\omega)\mathrm{d}\omega=\mathrm{d}F(\omega)$, and $f(\omega)$ forms a Fourier pair with $\gamma(\tau)$ so that
@@ 2.3 Welch’s method for spectral density estimation
where $S$ is an integer-valued shift factor satisfying $0<S\leq L$, so that $p=(L-S)/L$, where $10\leq p<$ 1, is the percentage overlap of the blocks, and the relationship between $M$, $L$ and $p$ is
$L[M(1-p)+p]=n,$ (8)
@@
where $\rho(\omega;|m-m^{\prime}|,h)=\mathrm{cor}\left[I_{L}^{m}(\omega;h),I_{L}^{m^{\prime}}(\omega;h)\right]$ and $\mathrm{var}[I_{L}(\omega;h)]=\mathrm{var}[I_{L}^{m}(\omega;h)]$ for all $m$. When $n>>M$, Welch’s estimator is approximately equal to a lag-window estimator where the lag-window is given by the convolution of $h_{t}$ with itself. This is most simple to see for Bartlett’s estimator ($p=0$ and $h_{t}=L^{-1/2}$) where the lag-window is a triangular function of width $2L$, which results from the discrete convolution of $h_{t}=L^{-1/2}$. The primary focus of this manuscript is to study the debiasing of Welch’s estimator; although, we note, that the presented results may also be readily extended to certain lag-window estimators. We now establish some important properties of Welch’s estimator in Theorem 1.
@@
Assume a fourth-order stationary stochastic process $\{X_{t}\}$ with finite fourth-order moments and absolutely convergent fourth-order cumulant function. Welch’s estimator, defined in (9), of the observed process $\mathrm{X}_{n}=(x_{0},\dots,x_{n-1})$, with data taper $\{h_{t}\}$ that satisfies $\sum h_{t}^{2}=1$, has the following properties.
@@
Proof.
Proofs are provided in the Supplementary Material. ∎
@@
As $n$ increases, then it is not possible to increase both $L$, the block length, and $M$, the number of blocks, linearly with $n$ for a fixed overlap $p$, see (8); a trade-off is required. For fixed $L$ and linearly increasing $M$, Theorem 1a establishes that Welch’s estimate has the optimal convergence rate of $\mathcal{O}_{p}(M^{-1/2})=\mathcal{O}_{p}(n^{-1/2})$, but the estimator is inconsistent for all processes except white noise due to the bias inherent in $f(\omega)*\mathcal{H}_{L}(\omega)$ for fixed $L$ as discussed. On the other extreme, for fixed $M$ and linearly increasing $L$, the bias asymptotically vanishes but the variance is $\mathcal{O}_{p}(1)$ (Theorem 1a) and therefore the Welch estimate suffers the same inconsistency as the periodogram (Theorem 1b). Consistency is only achieved by increasing both $L$ and $M$ with $n$ at reduced rates, say, $L=\mathcal{O}(n^{\alpha})$ and $M=\mathcal{O}(n^{1-\alpha})$ for $10<\alpha<$ 1, to satisfy (8). In practice, regularity assumptions may be employed to aid the selection of $L$ and $M$. For example, if we assume $h_{t}=L^{-1/2}$ and $f(\omega)$ to be Lipschitz continuous then, as shown in Fejér (1910),
$\mathrm{bias}\left[\bar{I}_{L}(\omega;h)\right]=\mathrm{E}[\bar{I}_{L}(\omega;h)]-f(\omega)=\mathcal{O}_{p}\left(\frac{\log L}{L}\right)$
and so the optimal value of $\alpha$, with respect to mean squared error, is approximately $\alpha=1/3$. Further regularity conditions and specification of tapers may yield even faster convergence of the bias: assumptions on the order of differentiability of $f(\omega)$ can further reduce the optimal choice of $\alpha$ (see, for example, Zhurbenko, 1980), and tapers may be chosen with associated properties, such as the family of Slepian tapers that maximally concentrate bandwidth energy (Slepian, 1978).
@@ 3.3 Model fitting
The motivation for (16) is as follows. From Theorem 1b, as $\lim_{L\rightarrow\infty}\mathrm{var}\left[\bar{I}_{L}(\omega;h)\right]=c\;\mathrm{var}\left[I_{L}(\omega;h)\right]$, over all $\omega$, we substitute $\mathrm{var}\left[I_{L}(\omega;h)\right]$ for $\mathrm{var}\left[\bar{I}_{L}(\omega;h)\right]$ in (15). Further, as we establish in Theorem 2 below, $\bar{I}_{L}(\omega;h)^{2}$ converges in $M$ and $L$ to $\mathrm{var}[I_{L}(\omega;h)]$, and so we approximate $\mathrm{var}[I_{L}(\omega;h)]$ by $\bar{I}_{L}(\omega;h)^{2}$, resulting in (16). From $\hat{\vartheta}=(\hat{a}_{1},\dots,\hat{a}_{K})$, we define a debiased functional estimate of $f(\omega)$ as $f(\omega;\hat{\vartheta})=\sum_{k=1}^{K}\hat{a}_{k}B_{k}(\omega)$. For the symmetric rectangular basis function of Section 3.2, we define the debiased Welch estimator at a discrete set of frequencies as
$\hat{I}(\omega_{k})=\hat{a}_{k}B_{k}(\omega_{k})=\hat{a}_{k}/\delta_{k},\quad\omega_{k}\in\{\omega_{1}^{\mathrm{c}},\dots,\omega_{K}^{\mathrm{c}}\};$
this is analogous to the frequencies for which Welch’s estimator is generally defined. Due to finite-sample error, the debiased Welch estimator is not guaranteed to yield a non-negative solution. We address this in Section 3.4 by constraining the solution space of $\hat{\vartheta}$ so that ϑ^+^italic-$\hat{\vartheta}\in\mathcal{R}^{+}$ where $\mathcal{R}^{+}$ denotes the non-negative space of reals. As by definition, $f(\omega)\in\mathcal{R}^{+}$, and following similar arguments to Theorem 2 of Politis and Romano (1995), we can establish that constraining ϑ^+^italic-$\hat{\vartheta}\in\mathcal{R}^{+}$ yields a better estimator with respect to mean squared error.
@@
Assume a fourth-order stationary process, $\{X_{t}\}$, with finite fourth-order moments and absolutely convergent fourth-order cumulant function, $\mathcal{K}(\tau_{i},\tau_{j},\tau_{k})$. The variance of the periodogram of any observed length-$L$ segment of $\mathrm{X}_{n}=(x_{0},\dots,x_{n-1})$, for all $\omega\neq 0,\omega_{\mathcal{N}}$, and with data taper $\{h_{t}\}$ that satisfies $\sum h_{t}^{2}=1$, is
@@
    $\mathrm{var}[I_{L}(\omega;h)]=\bar{I}_{L}(\omega;h)^{2}+\mathcal{O}_{p}\left(\frac{1}{\sqrt{M}}+\frac{\log^{2}L}{L^{2}}\right)$, when $\mathcal{K}(\tau_{i},\tau_{j},\tau_{k})=0$ for all $\tau_{i},\tau_{j},\tau_{k}$; and
@@
Proof.
The proof is provided in the Supplementary Material. ∎
@@
From Theorem 1 of Sykulski et al. (2019), we find that the debiased Welch estimator, $\hat{I}(\omega_{k})$, when computed through (6), is unbiased so that $\mathrm{E}[\hat{I}(\omega_{k})]=f(\omega_{k};\vartheta)$. The bias that results from the Riemannian basis approximation can be readily quantified. As $f(\omega)$ has a bounded first derivative $f^{\prime}(\omega)$, the bias that results from the Riemann approximation is $\mathrm{bias}[f(\omega;\vartheta)]=\mathcal{O}_{p}(L^{-1})$. Following Chui (1971), if we further assume that $f^{\prime}(\omega)$ is of bounded variation, we may strengthen this statement to say $\mathrm{bias}[f(\omega;\vartheta)]=\mathcal{O}_{p}(L^{-2})$. As $\hat{I}(\omega_{k})$ is a linear transform of $\bar{I}_{L}(\omega;h)$ we find that $\mathrm{var}[\hat{I}(\omega_{k})]=C(K,L,h)\mathrm{var}[\bar{I}_{L}(\omega;h)]=\mathcal{O}_{p}(M^{-1})$ where $C(K,L,h)>0$ is a constant that depends on the choice of $K$, $L$ and the data taper, and is not necessarily bounded above by 1. Thus, in terms of mean squared error, the optimal parametrisation is $L=\mathcal{O}(n^{1/3})$, or $L=\mathcal{O}(n^{1/5})$ if $f^{\prime}(\omega)$ is of bounded variation. The weighted least squares routine in (16) follows asymptotic arguments in Theorems 1 and 2 that result in an estimator that is asymptotically unbiased and $\mathcal{O}_{p}(M^{-1/2})$ as long as $L$ and $M$ both increase with $n$. As $C(K,L,h)$ is not bounded above by $1$, the debiased Welch estimator is not guaranteed to have a smaller mean squared error as compared to the Welch estimator given the same parameterisation. However, as the debiased Welch estimator corrects for the effects of bias, in practice we may apportion between $L=\mathcal{O}(n^{\alpha})$ and $M=\mathcal{O}(n^{1-\alpha})$ with a lower $\alpha$ than would be chosen for Welch’s estimator to minimise mean squared error (see also Remark 1), allowing for extra variance reduction, as we shall demonstrate for finite samples in Section 4.2. This is a similar result to Politis and Romano (1995) who study bias-correction of lag-window estimators.